I have one more word problem that I cannot do correctly. I need some serious hlp.
The hockey team receives 2 points when they win and 1 point when they tie. One season a team won the championship with 69 points. They won 12 more games than they tied. How many wins and how many ties did the team have?
Help please, I still do not get how to do these........

Answer by stanbon(75887)

The hockey team receives 2 points when they win and 1 point when they tie. One season a team won the championship with 69 points. They won 12 more games than they tied. How many wins and how many ties did the team have?
----------
Let # of wins be "w"; Let # of ties be "t"
---------------------
Games equation: w = t+12
--------------------------------
Points equation: 2w + t = 69
---------------------
Substitute w=t+12 into the 2nd equation to solve for "t":
2(t+12) + t = 69
2t+24 + t = 69
3t = 45
t = 15 (# of ties)
w = t+12 = 27 (# of wins)
--------------
Cheers,
Stan H.